OpenStudy (anonymous):

Create a unique example of dividing a polynomial by a monomial and provide the simplified form. Explain, in complete sentences, the two ways used to simplify this expression and how you would check your quotient for accuracy.

OpenStudy (anonymous):

@lilkittykat36 @dmezzullo @Hero @satellite73

OpenStudy (anonymous):

@dmezzullo lol your sense of humor is unique but I found that funny lol can you help with this though ?

OpenStudy (dmezzullo):

lol ty, unfortunately no)': i can call 911 they will help

OpenStudy (anonymous):

Do you know someone that Can atleast explain it clearly for me its for my virtual schools class @dmezzullo

OpenStudy (dmezzullo):

Well i actually can answer that with an answer other than when in doubt dial 911

OpenStudy (dmezzullo):

Or stop drop and roll

OpenStudy (anonymous):

IDU

OpenStudy (dmezzullo):

Ermm who can help u? OH OH OH nvm.. Ummmm

OpenStudy (dmezzullo):

idk sorry, @Hero SHould be able to help u or @Mertsj

OpenStudy (anonymous):

@dmezzullo thanks :) love the sense of humor keep it up lol

OpenStudy (mertsj):

Do you know what a polynomial is?

OpenStudy (anonymous):

@Mertsj i do know what it is but no clear understanding of it .

OpenStudy (mertsj):

This is a polynomial. You can make up one of your own if you wish: \[15x^3-9x^2+27x\]

OpenStudy (mertsj):

Here is a monomial: 3x

OpenStudy (mertsj):

Or you could make up one of your own.

OpenStudy (anonymous):

@Mertsj so 10x^2-9x^3+15x ?

OpenStudy (mertsj):

Yes but put the term with the highest exponent first.

OpenStudy (anonymous):

@Mertsj oh okay so what do i do next ?

OpenStudy (mertsj):

It said to divide by a monomial so you had better make up a monomial and then divide.

OpenStudy (anonymous):

@Mertsj okkay x+8?

OpenStudy (mertsj):

a monomial is a polynomial that has only 1 term.

OpenStudy (mertsj):

So maybe just x or just 8. They are both monomials.

OpenStudy (anonymous):

@Mertsj okay i think i got it but i still feel uttelry confused thanks though :)

OpenStudy (mertsj):

Did you do the division?

OpenStudy (mertsj):

\[\frac{-9x^3+10x^2+15x}{x}=\frac{-9x^3}{x}+\frac{10x^2}{x}+\frac{15x}{x}=-9x^2+10x+15\]

OpenStudy (mertsj):

Or you could write it like this:

OpenStudy (mertsj):